(1/x)-(3/y)=-8/3
(4/x)+(2/y)=-10/3
find x, & y.

Answer by Alan3354(69443)

(1/x)-(3/y)=-8/3
(4/x)+(2/y)=-10/3
find x, & y
---------------
Sub r for 1/x and s for 1/y
---
r - 3s = -8/3 --> r = 3s - 8/3
4r+ 2s = -10/3
4(3s - 8/3) + 2s = -10/3
12s - 32/3 + 2s = -10/3
14s = 22/3
s = 11/21
x = 21/11
-------
etc